# Learning novel representations of variable sources from multi-modal Gaia data via autoencoders

### v1

@@ 3.2 Data used for evaluation purposes
  • Eclipsing binary stars divided into 1736 Algol-type (ECLa), 2672 $\beta$ Lyrae-type (ECLb) and 1890 W Ursae Majoris-type (ECLw).
  • Long Period Variables (LPV) of various types with 10,751 sources in total.
@@ 4.1 VAE for XP mean spectra
After preprocessing, the data is then compressed and reconstructed using a VAE with an architecture based on fully connected (FC) layers that is showcased in Fig. 2. Three layers are used for the encoder and decoder networks, respectively. Intermediate layers have 128 neurons with Gaussian Error Linear Unit (GELU) activation functions (Hendrycks & Gimpel 2016) and layer normalisation (LN) to enhance stability during training (Lei Ba et al. 2016). The output layer of the encoder, i.e. the mean and standard deviation of the latent variables, do not incorporate LN and use a linear activation (LIN) to avoid constraining their range. The output layer of the decoder, i.e. the reconstructed mean spectra, uses the Rectified Linear Unit (ReLU) $\max(X,0)$ activation (Nair & Hinton 2010; Glorot et al. 2011) to ensure non-negativity. A separate two-layer decoder network is used to predict $\log\sigma_{\rm XP}^{2}$, an uncertainty term representing the logarithm of the variance of the reconstructed mean spectra. The particular choice for the number of neurons and activation functions of this network is detailed in Appendix B.
@@ 6 Identification of outliers using the latent variables
A widely used heuristic for anomaly detection that is specific to AEs consists of using the quality of the reconstructed data to identify outlier examples. This assumes that anomalous sources are harder to reconstruct because the model minimises the overall reconstruction error, thereby prioritising normal instances (Pang et al. 2020; Sánchez-Sáez et al. 2021). In this case, this corresponds to identifying sources for which the decoder variance $\sigma_{F}^{2}$ is large. In what follows, we use both the LOF and $\sigma_{F}^{2}$ to identify outlier candidates and analyse the differences between these approaches.
Figure 9 shows the LOF versus $\sigma_{F}^{2}$ distribution for the 7431 sources from Clementini et al. (2023) in the test-set partition of our training set. Most sources are concentrated around LOF values close to one and have low $\sigma_{F}^{2}$. In general, the correlation between LOF and $\sigma_{F}^{2}$ is weak, i.e. the most isolated sources are not necessarily due to the model being unable to reconstruct them correctly. The three sources with the highest LOF values are marked with stars. Notably, these sources also exhibit low values of $\sigma_{F}^{2}$, suggesting that their reconstructions are reliable. The folded light curves of these sources are shown in the left panel of Fig. 10, ranked by their LOF from top to bottom. The shape of the light curves suggests that these sources are likely eclipsing binaries. This is further confirmed using the amplitude ratio between the $G_{\rm BP}$ and $G_{\rm RP}$ band time series, as indicated by (Clementini et al. 2023). Moreover, the second and third sources have been classified as binaries using ATLAS data (Heinze et al. 2018). The full potential of latent-space neighbour analysis for identifying contaminants in Gaia data will be examined in greater detail in Wyrzykowski et al. (in prep.).
@@ 7 Supervised classification in the latent space
  • Most LPVs are correctly recovered, albeit with significant contamination from YSO, ELL, $\delta\,$Cep and t2 Cep stars. The sources classified as LPV are characterised by their strong red spectra.
  • ACV—CP, BE—GCAS, $\beta\,$Cep stars and SPB form a cluster. All these classes are characterised by their strong blue spectra.
@@
  • Clusters associated with RR Lyr ab and RR Lyr c stars due to their constrained period range.
  • A long time-scale cluster with the LPV, AGN, CV and a considerable fraction of $\delta\,$Cep and t2 Cep stars.
  • A cluster with the ECLa and a fraction of the ECLb, due to the model capturing the pattern of the top and bottom magnitude bins being filled by the sharp eclipses of these sources.
@@
  • Eclipsing Binaries (ECL) yield recalls above 50%, with most of the contamination confined to their sub-classes. This demonstrates the model’s ability to effectively capture the distinctive phase patterns of these sources.
@@
  • A cluster containing sources from SOLAR, YSO, SPB, $\beta\,$Cep and $\gamma\,$Dor stars, among other classes. The folded light curves of these sources exhibit shapes resembling noisy sine waves.
  • A cluster comprising LPVs, AGNs, CVs, YSO, and other classes. This grouping is primarily composed of sources that were folded with incorrectly estimated periods and/or exhibiting highly irregular behaviour.
Figure 11(d) shows the results of the LR classifier using the latent variables from all modalities as input. For reference, the precision, recall, and $F_{1}$-scores per class for this classifier are listed in Table 4. Performance is superior to the individual modalities across all individual classes, which is reflected in the strong diagonal of the confusion matrix. The classes that benefit less from the multimodal approach are LPV and WD, as these were already well separated using solely the XP spectra. On the other hand, the following classes exhibit a relative increase in $F_{1}$-score of more than 50% with respect to the best single modality: $\beta\,$Cep stars, $\gamma\,$Dor stars, ELL, RS, SPB, RR Lyr d stars, t2 Cep, CV and YSO.
@@
  • RS, ELL and YSO: Confused between them and with the SOLAR class.
  • SPB: This minor class has significant contamination from ACV—CP, BE—GCAS and $\beta\,$Cep stars due to their similar XP properties and periodicities.
  • RR Lyr d stars and ECLb are difficult to distinguish from other members of their parent class.
@@
This comparison is carried out using a subset of 3300 sources obtained by removing from our test set all the sources that Rimoldini et al. (2023) used for training. The results presented in Table 2 show overall agreement between the two, though significant differences arise in some of the classes, such as ACV, CEP, SOLAR and YSO. However, given the limited sample size, these differences may not be fully representative of the broader dataset. Notably, both models fail entirely for the RS class, highlighting the difficulty of accurately classifying these stars using Gaia data.
@@ 9 Astrophysical interpretation of the latent space
$\delta\,$Sct stars, $\gamma\,$Dor stars, and white dwarfs
@@
Eclipsing binary stars
The overdensities in the 2D latent space around $(\lambda_{1},\lambda_{2})=(-0.5,-1.4)$ are dominantly occupied by binary stars as shown in Fig. 14. The substructure in that region is largely determined by the morphology of the (folded) light curve. To demonstrate this, we follow Mowlavi et al. (2023) who used a simple but effective model to characterise the variability of Gaia DR3 binary candidates. In their work, the eclipses were modelled using two Gaussian functions, while an ellipsoidal variation was modelled with a cosine function. In Figs. 18, 34 and 35, we plot for the binary candidates in their catalogue, respectively, the magnitude difference between two consecutive eclipses, the duration of the primary eclipse as a fraction of the orbital period, and the amplitude of the cosine component (if any), which serves as a proxy for the amplitude of the ellipsoidal variation.
@@
Long-period variables
@@
Active galactic nuclei
@@ Acknowledgements.
PH, JDR and CA acknowledge support from the BELgian federal Science Policy Office (BELSPO) through various PROgramme de Développement d’Expériences scientifiques (PRODEX) grants. CA, PH, and MV acknowledge financial support from the European Research Council (ERC) under the Horizon Europe programme (Synergy Grant agreement N◦101071505: 4D-STAR). While partially funded by the European Union, views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. PH also acknowledges support from ANID – Millennium Science Initiative Program – ICN12_009 awarded to the Millennium Institute of Astrophysics MAS. CA, AK, DJF and HW acknowledge financial support from the Flemish Government under the long-term structural Methusalem funding program by means of the project SOUL: Stellar evOlution in fUll gLory, grant METH/24/012 at KU Leuven. VV gratefully acknowledges support from the Research Foundation Flanders (FWO) under grant agreement N°1156923N (PhD Fellowship). This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This work has made use of the Python package GaiaXPy, developed and maintained by members of the Gaia Data Processing and Analysis Consortium (DPAC) and in particular, Coordination Unit 5 (CU5), and the Data Processing Centre located at the Institute of Astronomy, Cambridge, UK (DPCI). This research was supported by the Patagón supercomputer of Universidad Austral de Chile (FONDEQUIP EQM180042).
@@ Appendix A Supporting figures and tables
Figure 21: Input data and VAE reconstructions for three Gaia DR3 sources belonging to the ACV, BCEP and SPB classes, respectively. The description of the rows and columns follows that of Fig. 1.
Figure 22: Input data and VAE reconstructions for three Gaia DR3 sources belonging to the CEPd, DSCT and GDOR classes, respectively. The description of the rows and columns follows that of Fig. 1.
@@
Figure 34: The same as for Fig. 18, but here the colour map represents the duration of the primary eclipse expressed as the fraction of the phase.
Figure 35: The same as for Fig. 18, but here the colour map represents the amplitude of the ellipsoidal variation (if any) in mmag.
Figure 36: The same as for Fig. 19, but here the main period $P_{1}$ found in the Lomb-Scargle periodogram was used for the colour coding.
@@ Appendix B Neural network architecture details
The architecture of the encoder ANN of the VAE for sampled XP low-resolution mean spectra is described as follows:
@@
The $\Delta m|\Delta t$ representation is compressed using an encoder ANN composed of FC layers. The architecture is described as follows:
@@
    A GRU [^8] Preliminary experiments showed that the difference in performance between GRU and the more complex Long Short Term Memory (Hochreiter et al. 1997) was negligible, hence the former being more efficient was preferred. with input size $(1+2H)$, hidden state size of $128$ and dropout probability of $0.2$, receives $E_{1}$ and produces a real-valued array $E_{2}$ with dimensionality $L\times 128$.
@@
    The phases are expanded using the same procedure described for the encoder, and concatenated with the latent variables forming a real-valued array $D_{1}$ with dimensionality $L\times(K+2H)$.

### v2

@@ 3.2 Data used for evaluation purposes
  • Eclipsing binary stars (ECL) divided into 1736 Algol-type (ECLa), 2672 $\beta$ Lyrae-type (ECLb) and 1890 W Ursae Majoris-type (ECLw).
  • Long Period Variables (LPV) including Mira variables and various semiregular variables with 10,751 sources in total.
@@ 4.1 VAE for XP mean spectra
After preprocessing, the data is then compressed and reconstructed using a VAE with an architecture based on fully connected (FC) layers that is showcased in Fig. 2. Three layers are used for the encoder and decoder networks, respectively. Intermediate layers have 128 neurons with Gaussian Error Linear Unit (GELU) activation functions (Hendrycks & Gimpel 2016) and layer normalisation (LN; Lei Ba et al. 2016) to enhance stability during training. The output layer of the encoder, i.e. the mean and standard deviation of the latent variables, do not incorporate LN and use a linear activation (LIN) to avoid constraining their range. The output layer of the decoder, i.e. the reconstructed mean spectra, uses the Rectified Linear Unit (ReLU) $\max(X,0)$ activation (Nair & Hinton 2010; Glorot et al. 2011) to ensure non-negativity. A separate two-layer decoder network is used to predict $\log\sigma_{\rm XP}^{2}$, an uncertainty term representing the logarithm of the variance of the reconstructed mean spectra. The particular choice for the number of neurons and activation functions of this network is detailed in Appendix B.
@@ 6 Identification of outliers using the latent variables
A widely used heuristic for anomaly detection that is specific to AEs consists of using the quality of the reconstructed data to identify outlier examples. This assumes that anomalous sources are harder to reconstruct because the model minimises the overall reconstruction error, thereby prioritising normal instances (Pang et al. 2021; Sánchez-Sáez et al. 2021). In this case, this corresponds to identifying sources for which the decoder variance $\sigma_{F}^{2}$ is large. In what follows, we use both the LOF and $\sigma_{F}^{2}$ to identify outlier candidates and analyse the differences between these approaches.
Figure 9 shows the LOF versus $\sigma_{F}^{2}$ distribution for the 7431 sources from Clementini et al. (2023) in the test-set partition of our training set. Most sources are concentrated around LOF values close to one and have low $\sigma_{F}^{2}$. In general, the correlation between LOF and $\sigma_{F}^{2}$ is weak, i.e. the most isolated sources are not necessarily due to the model being unable to reconstruct them correctly. The three sources with the highest LOF values are marked with stars. Notably, these sources also exhibit low values of $\sigma_{F}^{2}$, suggesting that their reconstructions are reliable. The folded light curves of these sources are shown in the left panel of Fig. 10, ranked by their LOF from top to bottom. The shape of the light curves suggests that these sources are likely eclipsing binaries. This is further confirmed using the amplitude ratio between the $G_{\rm BP}$ and $G_{\rm RP}$ band time series, as indicated by Clementini et al. (2023). Moreover, the second and third sources have been classified as binaries using ATLAS data (Heinze et al. 2018). The full potential of latent-space neighbour analysis for identifying contaminants in Gaia data will be examined in greater detail in Wyrzykowski et al. (in prep.).
@@ 7 Supervised classification in the latent space
  • Most LPVs are correctly recovered. The sources classified as LPV are characterised by their strong red spectra.
  • SPB stars are the primary source of misclassification for ACV—CP and $\beta\,$Cep stars. Conversely, the most significant source of confusion for the SPB stars are ACV—CP and $\beta\,$Cep stars. Additionally, $\beta\,$Cep stars are the leading source of misclassification for BE—GCAS. All these classes are characterised by their strong blue spectra.
@@
  • Most RR Lyr ab and RR Lyr c stars are correctly recovered due to their constrained period range. RR Lyr d stars, which pulsate in double mode are mostly confused with the former.
  • $\delta\,$Cep and t2 Cep stars are confused with one another, as well as with LPVs. The period range of the classes partially overlap. Furthermore, some AGN and CV are also misclassified as LPV. All of these sources are characterized by variability on long timescales.
  • Most ECLa are correctly recovered. This is due to the model capturing the pattern of the top and bottom magnitude bins being filled by the sharp eclipses of these sources.
@@
  • ECL yield recalls above 50%, with most of the contamination confined to their sub-classes. This demonstrates the model’s ability to effectively capture the distinctive phase patterns of these sources.
@@
[…]
  • AGNs are frequently misclassified as YSOs, WDs, and LPVs. Similarly, LPVs are often misclassified as YSOs and AGNs. For non-periodic and semiregular variables, the estimated periods are, in most cases, related to Gaia scan-angle dependent signals (Holl et al. 2023).
Figure 11(d) shows the results of the LR classifier using the latent variables from all modalities as input. For reference, the precision, recall, and $F_{1}$-scores per class for this classifier are listed in Table 4. Performance is superior to the individual modalities across all individual classes, which is reflected in the strong diagonal of the confusion matrix. The following classes exhibit a relative increase in $F_{1}$-score of more than 50% with respect to the best single modality: $\beta\,$Cep stars, $\gamma\,$Dor stars, ELL, RS, SPB, RR Lyr d stars, t2 Cep, CV and YSO.
@@
  • RS, ELL and YSO: These classes are confused between them and with the SOLAR class. RS, ELL and SOLAR are all examples of rotational variables. Confusion between these classes was also reported by Rimoldini et al. (2023) (cf. Table 3).
[…]
  • RR Lyr d stars, which are multimode pulsators, are difficult to distinguish from other members of their parent class. A similar situation can be observed for ECLb. The folded light curve morphology of ECLb is a transition between ECLa and ECLw and this is captured by the latent variables.
@@
This comparison is carried out using a subset of 3300 sources obtained by removing from our test set all the sources that Rimoldini et al. (2023) used for training. The results presented in Table 2 show overall agreement between the two, though significant differences arise in some of the classes, such as ACV, CEP, SOLAR and YSO. However, given the limited sample size, these differences may not be fully representative of the broader dataset. Notably, both models are not able to recover the RS sources in this sample.
@@ 9.1 RR Lyr stars
9.2 $\delta\,$Sct stars, $\gamma\,$Dor stars, and white dwarfs
@@ 9.3 Hot variable stars
9.4 Eclipsing binary stars
The overdensities in the 2D latent space around $(\lambda_{1},\lambda_{2})=(-0.5,-1.4)$ are dominantly occupied by binary stars as shown in Fig. 14. The substructure in that region is largely determined by the morphology of the (folded) light curve. To demonstrate this, we follow Mowlavi et al. (2023) who used a simple but effective model to characterise the variability of Gaia DR3 binary candidates. In their work, the eclipses were modelled using two Gaussian functions, while an ellipsoidal variation was modelled with a cosine function. In Figs. 18, 34 and 36, we plot for the binary candidates in their catalogue, respectively, the magnitude difference between two consecutive eclipses, the duration of the primary eclipse as a fraction of the orbital period, and the amplitude of the cosine component (if any), which serves as a proxy for the amplitude of the ellipsoidal variation.
@@ 9.4 Eclipsing binary stars
9.5 Long-period variables
@@
9.6 Active galactic nuclei
@@ Acknowledgements.
PH, JDR and CA acknowledge support from the BELgian federal Science Policy Office (BELSPO) through various PROgramme de Développement d’Expériences scientifiques (PRODEX) grants. JDR and CA acknowledge funding from the Research Foundation Flanders (FWO) under grant agreement G089422N. CA, PH, and MV acknowledge financial support from the European Research Council (ERC) under the Horizon Europe programme (Synergy Grant agreement N◦101071505: 4D-STAR). While partially funded by the European Union, views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. PH also acknowledges support from ANID – Millennium Science Initiative Program – ICN12_009 awarded to the Millennium Institute of Astrophysics MAS. CA, AK, DJF and HW acknowledge financial support from the Flemish Government under the long-term structural Methusalem funding program by means of the project SOUL: Stellar evOlution in fUll gLory, grant METH/24/012 at KU Leuven. VV gratefully acknowledges support from the Research Foundation Flanders (FWO) under grant agreement N°1156923N (PhD Fellowship). This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This work has made use of the Python package GaiaXPy, developed and maintained by members of the Gaia Data Processing and Analysis Consortium (DPAC) and in particular, Coordination Unit 5 (CU5), and the Data Processing Centre located at the Institute of Astronomy, Cambridge, UK (DPCI). This research was supported by the Patagón supercomputer of Universidad Austral de Chile (FONDEQUIP EQM180042).
@@ Appendix A Supporting figures and tables
Figure 21: Input data and VAE reconstructions for three Gaia DR3 sources belonging to the ACV, $\beta\,$Cep and SPB classes, respectively. The description of the rows and columns follows that of Fig. 1.
Figure 22: Input data and VAE reconstructions for three Gaia DR3 sources belonging to the $\delta\,$Cep, $\delta\,$Sct and $\gamma\,$Dor star classes, respectively. The description of the rows and columns follows that of Fig. 1.
@@
Figure 34: The same as Fig. 18, but colour represents the duration of the primary eclipse expressed as the fraction of the phase.
Figure 35: The same as for Fig. 19, but the Lomb-Scargle dominant period $P_{1}$ is used for the colour coding.
Figure 36: The same as Fig. 18, but colour represents the amplitude of the ellipsoidal variation (if any) in mmag.
@@ Appendix B Neural network architecture details
The architecture of the encoder of the VAE for sampled XP low-resolution mean spectra is described as follows:
@@
The $\Delta m|\Delta t$ representation is compressed using an encoder composed of FC layers. The architecture is described as follows:
@@
The latent variables are sampled using Eq. (5) and given as input to a decoder with the following architecture:
@@
    A GRU [^8] Preliminary experiments showed that the difference in performance between GRU and the more complex LSTM (Hochreiter et al. 1997) was negligible, hence the former being more efficient was preferred. with input size $(1+2H)$, hidden state size of $128$ and dropout probability of $0.2$, receives $E_{1}$ and produces a real-valued array $E_{2}$ with dimensionality $L\times 128$.
@@
    The phases are expanded using the same procedure described for the encoder, and concatenated with the latent variables forming an array $D_{1}$ with dimensionality $L\times(K+2H)$.